1.If cos(b-c).cos(c-a).cos(a-b) + 1 = 0.


the prove that at least one of the angles (a-b),(b-c) and (c-a) is and odd multiple of ∏


2.A + B + C = 180 prove that cos A/2 + cos B/2 + cos C/2 = 4cos(∏-A)/4 . cos(∏-B)/4 . cos(∏-C)/4

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
13 years ago

Dear manoj

cos(b-c).cos(c-a).cos(a-b) + 1 = 0.

or   cos(b-c).cos(c-a).cos(a-b) =-1

only possible if

 cos(b-c)=cos(c-a)=cos(a-b) =-1

or

  cos(b-c)=cos(c-a)=1   and  cos(a-b) =-1

  cos(b-c)=1   , cos(c-a)=-1  and  cos(a-b) =1

  cos(b-c)=-1   ,  cos(c-a)=  cos(a-b) =1

in any case atleast one of three term will be eqaul to -1

 let cos(b-c) =-1

             b-c = (2n+1)

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin

Now you can win exciting gifts by answering the questions on Discussion Forum. So help

discuss any query on askiitians forum and become an Elite Expert League askiitian.